Part way through the production process it is discovered that the cost assignment was inaccurate for product A and that there has been an underestimation of $150 per unit. They are unable to change the pricing because of contractual obligations.

What is the impact of this error in cost assignment if sales were as predicted?

calculate increase/decrease of predicted sales using data given for product A 12000 units.

Predicted: 12,000 of product A and 8,000 of product B.
Actual: 15,000 for product A and 7,000 for product B.
The wholesale margin on product A is calculated to be $450 and on product B it is $350.
The predicted sales mix is 12,000/20,000 = 60% of A to 40% of B.
The actual sales mix is 15,000/22,000 = 68.2% of A to 31.8% of B.
The sales mix variance for A is 15,000*(0.682% - 0.6%)*$450 = 553,500 or favorable variance.
For B, the sales mix variance is 7,000*(0.318 - 0.4)*350 = -200,900 or unfavorable variance
1
Expert's answer
2018-08-29T12:49:09-0400
Predicted:
12 000*450 + 8 000*350 = 8 200 000.
Actual:
15 000*450 + 7 000*350 = 9 200 000,
this indicator does not take into account an underestimation of $ 150.
Then
15 000*(450-150) + 7 000*350 = 6 950 000.

8 200 000 – 6 950 000 = 1 250 000

Thus, the forecasted result for 1 250 000 more than the final result.
In the absence of underestimation, the actual result would exceed forecasts by
9 200 000 – 8 200 000 = 1 000 000.